# On the optimality of coin-betting for mean estimation

### v2

@@ 7 Perspectives
Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both extending the results on characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean as a special case. Their analyses, however, rely on considerably more abstract and technically advanced proof techniques. Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered. Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3. We also remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025). That said, for the multi-round setting, the approach does give rise to non-trivial measurability issues when $\mathcal{X}$ is uncountable. These difficulties already emerge in the two-round case: although one can show that every e-variable must be dominated by a product of the form $(1+\lambda_{1}x_{1})(1+\lambda_{2}(x_{1})x_{2})$, this is constructed in a way that gives no guarantee on the measurability of $\lambda_{2}$. Whether this issue can be resolved remains an open question. Nonetheless, even in the simplified setting we adopt by requiring $\mathcal{X}$ to have finite cardinality [^11] A countable $\mathcal{X}$ would also suffice to solve the measurability issue, though the final step of the proof of Theorem 3 would require more delicate handling., this work is, to the best of the author’s knowledge, the first to fully characterise all e-processes for the natural problem of testing a fixed conditional mean.
To conclude, let us remark that this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023). Finally, although we have focused on random variables taking values in $[0,1]$ for clarity of exposition, all the results presented readily extend to any bounded closed interval.
@@ Acknowledgements
The author would like to thank Gergely Neu, Peter D. Grünwald, Nishant A. Mehta, Hamish E. Flynn, and Claudia M. Chanu, for the insightful discussions that inspired this work. I would also like to thank Nick W. Koning and Aditya Ramdas for their constructive feedback, which helped improve this work following its initial appearance, and Wouter M. Koolen for helpful discussions on e-processes. The author acknowledges the use of GPT-4o during the redaction of this paper, in order to polish the presentation. All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
Appendix
Proof of Theorem 2
Proof.
We start by showing that the class of functions considered is an e-class, and that all its elements are maximal e-variables. So, fix any $\lambda^{T}_{\mu}\in\Lambda^{T}$. As all the $\lambda_{t}$ are valued in $I_{\mu}$, it is clear that $E_{\lambda^{T}}$ is non-negative. Moreover, for any $P\in\mathcal{H}_{\mu}^{T}$ it is easily checked that $\mathbb{E}_{P}[E_{\lambda^{T}}]=1$, which ensures that $E_{\lambda^{T}}$ is an e-variable. For the maximality we can use the same argument as in the proof of Theorem 1. Let $\hat{E}\in\mathcal{E}^{T}_{\mu}$ be such that $\hat{E}\succeq E_{\lambda^{T}}$. Fix any $x^{T}\in\mathcal{X}^{T}$. By Lemma 4, for each $t$ we can find a $P_{t}\in\mathcal{H}^{1}_{\mu}$ such that $P_{t}(\{x_{t}\})>0$. Let $P=P_{1}\otimes\dots\otimes P_{T}$. Then we have that $P(\{x^{T}\})>0$. In particular, we have that $0\leq P(\{x^{T}\})(\hat{E}(x^{T})-E_{\lambda^{T}}(x^{T}))\leq\mathbb{E}_{P}[\hat{E}-E_{\lambda^{T}}]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$. Since $P(\{x^{T}\})>0$, $\hat{E}(x^{T})=E_{\lambda^{T}}(x^{T})$, and we conclude that $E_{\lambda^{T}}$ is maximal.
To show that the e-class that we consider is majorising we proceed by induction on $T$. From Theorem 1 we know that the statement is true for $T=1$. For $T\geq 2$. let us assume that the claim holds for $T-1$, and let us show that it is also true for $T$. Let $E$ be an e-variable for $\mathcal{H}^{T}_{\mu}$. First, we show that, for any $Q\in\mathcal{H}^{T-1}_{\mu}$ and any $x\in\mathcal{X}$, we have
$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq F_{\mu}(x)\,,$
with $F_{\mu}$ as in Lemma 4. Fix $Q$ and $x\geq\mu$ in $\mathcal{X}$. Define $V_{x}=P_{x}\otimes Q$, where $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Clearly, $V_{x}\in\mathcal{H}_{\mu}^{T}$. We have that $1\geq\mathbb{E}_{V_{x}}[E]=\tfrac{\mu}{x}\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]+\big{(}1-\tfrac{\mu}{x}\big{)}\mathbb{E}_{Q}[E(0,X_{2},\dots,X_{T})]$, and so $\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq x/\mu=F_{\mu}(x)$. The case for $x\leq\mu$ is essentially analogous. In particular, the following two sets are non-empty:
$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu)\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,;$
$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1]\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,.$
$B_{0}$ and $B_{1}$ are intersections of closed intervals (one per each allowed $x$ and $Q$), and as such they must be closed intervals. We can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$. We will show that $\beta_{0}\geq\beta_{1}$, and hence that $B_{0}\cap B_{1}\neq\varnothing$. The argument is essentially the same used in the proof of Theorem 1 (cf. also Figure 1). Assume that $\beta_{0}<\beta_{1}$, and let $\beta^{\star}\in(\beta_{0},\beta_{1})$. As $\beta^{\star}\notin B_{0}$, there is $Q_{0}\in\mathcal{H}^{T-1}_{\mu}$ and $u_{0}<\mu$ in $\mathcal{X}$ such that $\mathbb{E}_{Q_{0}}[E(u_{0},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{0}-\mu)$. Similarly, there must be $Q_{1}\in\mathcal{H}_{\mu}^{T-1}$ and $u_{1}>\mu$ in $\mathcal{X}$, such that $\mathbb{E}_{Q_{1}}[E(u_{1},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{1}-\mu)$, since $\beta^{\star}\notin B_{1}$. Let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$. Then, $V\in\mathcal{H}^{T}_{\mu}$. Moreover,
which is a contradiction since $E\in\mathcal{E}^{T}_{\mu}$. We have thus established that $B_{0}\cap B_{1}$ is non-empty, and in particular there is $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$ and every $Q\in\mathcal{H}_{\mu}^{T-1}$,
$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\lambda_{1}(x-\mu)\,.$
Fix $y\in\mathcal{X}$ such that $1+\lambda_{1}(y-\mu)\neq 0$. Define $E_{y}:\mathcal{X}^{T-1}\to\mathbb{R}$ via $E_{y}(x^{2:T})=E(y,x^{2:T})/(1+\lambda_{1}(y-\mu))$, where we recall that $x^{2:T}$ is a shorthand for $(x_{2},\dots,x_{T})$. $E_{y}$ is non-negative, since by definition of $I_{\mu}$ we know that $1+\lambda(y-\mu)$ is non-negative for any $\lambda\in I_{\mu}$. From what we have already shown, $E_{y}$ is an e-variable for $\mathcal{H}^{T-1}_{\mu}$. Thanks to the inductive hypothesis, for each $t\in[2:T]$ there exist a function $\tilde{\lambda}_{t}^{y}:\mathcal{X}^{t-2}\to I_{\mu}$ such that $E_{y}(x^{2:T})\leq\prod_{t=2}^{T}(1+\tilde{\lambda}_{t}^{y}(x^{2:(t-1)}))(x_{t}-\mu))$, for any $x^{2:T}\in\mathcal{X}^{T-1}$. For all $t\in[2:T]$, and every $x^{t-1}\in\mathcal{X}^{t-1}$, define
$\lambda_{t}(x^{t-1})=\begin{cases}0&\text{if $1+\lambda_{1}(x_{1}-\mu)=0$};\\ \big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{t}^{x_{1}}(x^{2:(t-1)})&\text{otherwise.}\end{cases}$
So far, we have proved that for any $x^{T}\in\mathcal{X}^{T}$, such that $1+\lambda_{1}(x_{1}-\mu)\neq 0$, we have that $E(x^{T})\leq E_{\lambda^{T}}(x^{T})$. We only need to check that this is still true when $01+\lambda_{1}(x_{1}-\mu)=$ 0. In such case, $E$ is $Q$-almost surely null for all $Q\in\mathcal{H}^{T-1}_{\mu}$, which implies that $E(x^{T})=0$, and so we conclude. ∎
Proof of Theorem 3
Proof.
We start by recalling and introducing some definitions. For $T\geq 1$, $\Lambda^{T}_{\mu}$ is the set whose elements are $T$-tuple $\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})$, with $\lambda^{t}:\mathcal{X}^{t-1}\to I_{\mu}$. We endow $\Lambda^{T}_{\mu}$ with the product topology of point-wise convergence. Specifically, a sequence $(\lambda^{t})_{t\geq 0}\to\tilde{\lambda}^{t}$ if, and only if, $\lambda_{t}\to\tilde{\lambda}_{t}$ point-wise for all $t\leq T$. For convenience, we will henceforth refer to this topology as the $\Lambda$-topology. It is straightforward to check that under the $\Lambda$-topology $\Lambda_{\mu}^{T}$ is compact, since $I_{\mu}$ is bounded and $\mathcal{X}$ has finite cardinality. Our proof strategy will rely on an inductive reasoning. For this, it is convenient to introduce the notion of a $T$-e-process, for $T\geq 1$. This is a sequence $(E_{t})_{t\geq 0}$ of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that $\mathbb{E}_{P}[E_{\tau}]\leq 1$ for any $\tau\in\mathcal{T}_{T}$, and any $P\in\mathcal{H}_{\mu}^{\infty}$. We remark that an e-process is a $T$-e-process for all $T\geq 1$.
We are now ready to start the proof. First, we will show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class. For any $T\geq 1$ and any $T$-e-process $E$, let $\tilde{\Lambda}^{T}_{E,\mu}\subseteq\Lambda^{T}_{\mu}$ denote the set of the $\lambda^{T}$ such that $E_{t}\preceq E_{\lambda^{t}}$, for all $t\in[1:T]$. We note that this is a closed and compact set under the $\Lambda$-topology that we introduced earlier. We now show that for every $T\geq 1$ and every $T$-e-process $E$, the set $\tilde{\Lambda}^{T}_{E,\mu}$ is non-empty. We prove this claim by induction. Since $E_{1}$ must be an e-variable for $\mathcal{H}_{\mu}^{1}$, as a consequence of Theorem 1, $\tilde{\Lambda}^{1}_{E,\mu}$ is non-empty. Now, fix $T\geq 2$ and a $T$-e-process $E$, and assume that, for any $(T-1)$-e-process $E^{\prime}$, $\tilde{\Lambda}^{T-1}_{E^{\prime},\mu}$ is non-empty. We will show that this implies that $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty. Fix an arbitrary $x\in\mathcal{X}$ and define the process $E^{x}=(E^{x}_{t})_{t\geq 1}$ such that $E_{t}^{x}(x^{t})=E_{t+1}(x,x_{1},\dots,x_{t})$, for any $x^{t}\in\mathcal{X}^{t}$. Define the sets
$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu),\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,;$
$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1],\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,.$
We claim that $B_{0}$ and $B_{1}$ are non-empty. Indeed, we now show that $\mu^{-1}\in B_{1}$. Analogously one can prove that $(\mu-1)^{-1}\in B_{0}$. The argument here follows closely the one depicted in Figure 1. Fix $x\in\mathcal{X}\cap[0,\mu)$, $Q\in\mathcal{H}_{\mu}^{\infty}$, and $\tau\in\mathcal{T}_{T-1}$. Define $\tau_{x}:\mathcal{X}^{\infty}\to\mathbb{N}$ as $\tau_{x}(x^{\infty})=1+\tau(x^{2:\infty})$, if $x_{1}=x$, otherwise $\tau_{x}(x^{\infty})=1$. It is easily checked that $\tau\in\mathcal{T}_{T}$. Also, we let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x}\delta_{0})$, and we define $V_{x}=P_{x}\otimes Q$ [^12] Note that both $V_{x}$ and $Q$ are measures on $\mathcal{P}_{\mathcal{X}^{\infty}}$. Here we mean that $V_{x}(\{x^{\infty}\})=P_{x}(\{x_{1}\})Q(\{x^{2:\infty}\})$ for all $x^{\infty}\in\mathcal{X}^{\infty}$. Clearly, $V_{x}\in\mathcal{H}_{\mu}^{\infty}$. Since $E$ is a $T$-e-process, we have
$1\geq\mathbb{E}_{V_{x}}[E_{\tau_{x}}]=(1-\tfrac{\mu}{x})E_{1}(0)+\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\geq\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\,.$
So, $E_{Q}[E^{x}_{\tau}]\leq\frac{x}{\mu}=1+\mu^{-1}(x-\mu)$, and so $\mu^{-1}\in B_{1}$.
Since both $B_{0}$ and $B_{1}$ can be written as intersections of non-empty closed intervals, we can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$. As usual, we want to show that $B_{0}\cap B_{1}$ is non-empty, or equivalently that $\beta_{1}\leq\beta_{0}$. Assume that this was not the case and there is $\beta^{\star}\in(\beta_{0},\beta_{1})$. Since $\beta^{\star}\notin B_{0}$, there must be $u_{0}<\mu$ in $\mathcal{X}$, $Q_{0}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{0}\in\mathcal{T}_{T-1}$ such that $\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]>1+\beta^{\star}(u_{0}-\mu)$. Similarly, as $\beta^{\star}\notin B_{1}$, one can find $u_{1}>\mu$ in $\mathcal{X}$, $Q_{1}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{1}\in\mathcal{T}_{T-1}$, such that $\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1+\beta^{\star}(u_{1}-\mu)$. Define $\tau:\mathcal{X}^{\infty}\to\mathcal{N}$ via $\tau(x^{\infty})=1+\tau_{0}(x^{2:\infty})$, if $x_{1}<\mu$, $\tau(x^{\infty})=1+\tau_{1}(x^{2:\infty})$ if $x_{1}>\mu$, and $\tau(x^{\infty})=1$ if $x_{1}=\mu$. Also, let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$. Then, we have that
$\mathbb{E}_{V}[E_{\tau}]=\frac{u_{1}-\mu}{u_{1}-u_{0}}\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{1}}(X_{2},\dots,X_{T})]+\frac{\mu-u_{0}}{u_{1}-u_{0}}\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}(X_{2},\dots,X_{T})]>1\,.$
However, this is a contradiction since $\tau\in\mathcal{T}_{T}$, $V\in\mathcal{H}_{\mu}^{\infty}$, and $E$ is a $T$-e-process. We can thus conclude that $B_{0}\cap B_{1}\neq\varnothing$ and, in particular, there must be $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$, every $Q\in\mathcal{H}_{\mu}^{\infty}$, and every $\tau\in\mathcal{T}_{T-1}$, we have
$\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\lambda_{1}(x-\mu)\,.$
Now, for any arbitrary $x\in\mathcal{X}$ such that $1+\lambda_{1}(x-\mu)\neq 0$. Define the process $\tilde{E}^{(x)}=(\tilde{E}^{(x)}_{t})_{t\geq 0}$ via
$\tilde{E}^{(x)}_{t}(x^{t})=\frac{E_{t+1}(x,x_{1},\dots,x_{t})}{1+\lambda_{1}(x-\mu)}\,.$
Let us show that $\tilde{E}^{(x)}$ is a $(T-1)$-e-process. It is clearly non-negative, as $\lambda_{1}\in I_{\mu}$ implies $1+\lambda_{1}(x-\mu)\geq 0$. Fix $Q\in\mathcal{H}_{\mu}^{\infty}$ and $\tau\in\mathcal{T}_{T-1}$. From what we have shown, we have
$\mathbb{E}_{Q}[\tilde{E}^{(x)}_{\tau}]=\frac{\mathbb{E}_{Q}[E^{x}_{\tau}]}{1+\lambda_{1}(x-\mu)}\leq 1\,.$
In particular, the inductive argument yields the existence of $\tilde{\lambda}_{(x)}^{T-1}\in\Lambda^{T-1}_{\mu}$, such that for all $t\in[1:T-1]$ we have $\tilde{E}^{(x)}\preceq E_{\tilde{\lambda}^{t}_{(x)}}$. Let $\lambda^{T}\in\Lambda^{T}_{\mu}$ be defined via $\lambda_{t}(x^{T})=\big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{(x_{1}),t}(x_{2},\dots,x_{T})$, where we set $\tilde{\lambda}_{(x),t}$ identically equal to zero whenever $01+\lambda_{1}(x-\mu)=$ 0. We will now show that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$. Fix any $x^{T}\in\mathcal{X}^{T}$. If $1+\lambda_{1}(x_{1}-\mu)\neq 0$, then for any $t\in[1:T]$ we have that
$E_{t}(x^{t})=\big{(}1+\lambda_{1}(x_{t}-\mu)\big{)}E^{(x_{1})}_{t-1}(x_{2},\dots,x_{t})\leq\prod_{s=1}^{t}\big{(}1+\lambda_{s}(x^{s-1})(x_{s}-\mu)\big{)}=E_{\lambda^{t}}(x^{t})\,.$
If $01+\lambda_{1}(x_{1}-\mu)=$ 0, then, for every $t\in[1:T-1]$, $\mathbb{E}_{Q}[E_{t}^{x_{1}}]=0$ for every $Q\in\mathcal{H}^{\infty}$, which implies that $E_{t}^{x_{1}}$ is identically null. In particular, $E_{t}(x^{t})=0=E_{\lambda^{t}}(x^{t})$. This shows that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$, and so $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
What we have shown so far allows to say that given an e-process $E$, for every $T\geq 1$ we can find a $\lambda^{T}\in\Lambda^{T}$ such that $E_{t}\leq E_{\lambda^{t}}$ for every $t\in[1:T]$. However, this does not tell us yet that there is a $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$ such that the same property holds. To show this, we leverage the compactness of $\tilde{\Lambda}^{T}_{E,\mu}$ under the $\Lambda$-topology. For $T^{\prime}\geq T$, let $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ denote the set of $\lambda^{T}\in\Lambda^{T}$ such that there is $\tilde{\lambda}^{T^{\prime}}\in\tilde{\Lambda}_{E,\mu}^{T^{\prime}}$ satisfying, for all $t\in[1:T]$, $\lambda_{t}=\tilde{\lambda}_{t}$. It is easily checked that $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ is compact in the $\Lambda$-topology. Moreover, if $T^{\prime\prime}\geq T^{\prime}$, then $\tilde{\Lambda}^{T|T^{\prime\prime}}_{E,\mu}\subseteq\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$. Since the intersection of countably many non-empty compact sets is non-empty, we can define $\tilde{\Lambda}^{1|\infty}_{E,\mu}=\bigcap_{T\geq 1}\tilde{\Lambda}^{1|T}_{E,\mu}$, which is a non-empty set. Now, fix $\lambda_{1}\in\tilde{\Lambda}^{1|\infty}_{E,\mu}$. We have that $E_{1}\preceq E_{\lambda_{1}}$. By construction $\lambda_{1}\in\tilde{\Lambda}^{1|2}_{E,\mu}$, and so we can find a $\lambda_{2}$ such that $(\lambda_{1},\lambda_{2})\in\tilde{\Lambda}^{2}_{E,\mu}$. We can go on like this indefinitely. In particular, we can construct a sequence $\lambda^{\infty}$ such that $\lambda^{T}\in\tilde{\Lambda}^{T}_{E,\mu}$ for all $T\geq 1$. This concludes the proof that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
We are left with checking that every e-process in $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is maximal. Fix $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$, and let $E$ be an e-process such that $E\succeq E^{\lambda^{\infty}}$. Clearly, we must have $E_{0}=1=E_{0}^{\lambda^{\infty}}$. Now, fix $T\geq 1$. We have that $E_{T}\succeq E_{\lambda^{T}}$. Since $E$ and $E^{\lambda^{\infty}}$ are e-processes, choosing the stopping time that is identically equal to $T$ shows that $E_{T}$ and $E_{\lambda^{T}}$ are $T$-round e-variables. In particular, by Theorem 2, $E_{\lambda^{T}}$ is maximal, and so $E_{T}\succeq E_{\lambda^{T}}$ implies that $E_{T}$ and $E_{\lambda^{T}}$ coincide. Since this is true for all $T\geq 1$, we conclude. ∎
Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$

### v4

@@ 7 Perspectives
Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean via single-round e-variables as a special case. Their results directly imply our Theorem 1 (but not Theorems 2 and 3) [^18] Interestingly, one could see the multi-round hypothesis $\mathcal{H}_{\mu}^{T}$ as a linearly constrained hypothesis, in the sense of Larsson et al. (2025). However, it does not seem to be trivial to directly derive Theorem 2 from their results.. However, their analyses rely on considerably more abstract and technically advanced proof techniques. Clerico (2024) first proves results for finite domains and then extends them to the uncountable case by compactness and density arguments, while Larsson et al. (2024) leverages and extends powerful duality tools from the theory of functional lattices. Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered. Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3. We remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025).
Another interesting connection with this paper’s approach and the literature is recent work on admissible merging for e-variables (Vovk and Wang, 2024; Wang, 2025). A thorough exploration of these connections is an interesting avenue for future research. In particular, Wang (2025) implies that the coin-betting single-round e-variables cannot be majorised by any e-variable that is a monotonic function. However, the current paper approach relies on this monotonicity assumption, and cannot hence imply directly Theorem 1. On the other hand, Vovk and Wang (2024) considered conditional sequential hypotheses, of which conditional mean testing is a special case. Although their framework adopts a slightly different perspective, some of their results may imply, or be equivalent to, a weaker version of Theorem 3, where a finite time horizon $T$ is fixed and the e-processes are defined as vectors $(E_{t})_{t\in[0:T]}$, rather than sequences. As our proof techniques follow a different route compared to these two works, we believe our findings offer an alternative perspective that contribute to a broader understanding of the problem and could potentially help generalise or strengthen their results.
To conclude, we remark that the present paper does not aim to delve into the design of effective coin-betting strategies. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023). However, let us stress once more that characterising the optimal e-class simplifies the taks of designing such strageties, by provides the minimal set where e-values shall be picked. We discussed at the end of Section 5 an explicit example where this can be turned into a practical advantage, namely when using a Bayesian aggregation strategy that places a prior over the e-class considered. The tractable, parametric form of the coin-betting e-class makes such approach feasible, while its optimality ensures that no statistical power is lost.
As a final comment, although we have focused on random variables taking values in $\mathcal{X}\subseteq[0,1]$ for clarity of exposition, most results presented can be easily extended to any bounded closed real set.
@@ Acknowledgements
The author would like to thank Gergely Neu, Peter D. Grünwald, Nishant A. Mehta, Hamish E. Flynn, and Claudia M. Chanu, for the insightful discussions that inspired this work. I would also like to thank Nick W. Koning and Aditya Ramdas for their constructive feedback, which helped improve this work following its initial appearance, and Wouter M. Koolen for helpful discussions on e-processes. Finally, I thank the anonymous reviewers for the insightful comments and feedback that helped improving this work. GPT-4o was used during the redaction of this paper to polish the presentation. All AI-generated text was reviewed and edited by the author, who takes full responsibility for the content of this manuscript. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
Appendix A
A.1 Proof of Proposition 2
Proof 7
For the first claim, without loss of generality we let $\mathcal{X}=[0,1]$, which implis the desired result for any $\mathcal{X}$. For any $\alpha\in\mathbb{R}$, let $S_{\alpha}:x\mapsto E^{\mathrm{H}}_{\alpha}(0)+(E^{\mathrm{H}}_{\alpha}(1)-E^{\mathrm{H}}_{\alpha}(0))x$ be the straight line that intersects $E^{\mathrm{H}}_{\alpha}$ at $x=0$ and $x=1$. Since $E^{\mathrm{H}}_{\alpha}$ is a convex function on [0,1][0,1][ 0 , 1 ], we have $E^{\mathrm{H}}_{\alpha}(x)\leq S_{\alpha}(x)$ for all $x\in[0,1]$. Let us show that $E_{\lambda_{\alpha}}(x)\geq S_{\alpha}(x)$ for all $x$. As $E_{\lambda_{\alpha}}$ and $S_{\alpha}$ represents parallel straight lines, it suffices to show inequality for a single point, say $x=\mu$ where $E_{\lambda_{\alpha}}$ equals 111. Hence, it is enough to prove that $1+(e^{\alpha}-1)\mu\leq e^{\alpha\mu+\alpha^{2}/8}$. As $\mu\in(0,1)$, we always have $1+(e^{\alpha}-1)\mu>0$, and we can define $u:\mathbb{R}\to\mathbb{R}$ as
$u(\alpha)=\frac{\alpha^{2}}{8}+\mu\alpha-\log\big{(}1+\mu(e^{\alpha}-1)\big{)}\,.$
Clearly, if $u$ is non-negative, then $1+(e^{\alpha}-1)\mu\leq e^{\alpha\mu+\alpha^{2}/8}$ for all $\alpha\in\mathbb{R}$, and so we obtain the desired claim. First, note that $u$ is twice differentiable, and we can explicitly compute its first and second derivatives:
$u^{\prime}(\alpha)=\frac{\alpha}{4}+\mu-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\quad\text{ and }\quad u^{\prime\prime}(\alpha)=\frac{1}{4}-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\left(1-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\right)\,.$
For any $\xi\in\mathbb{R}$ we have $\xi(1-\xi)\leq 1/4$, so $u^{\prime\prime}(\alpha)\geq 0$ for all $\alpha$, and $u$ is convex. Moreover $u^{\prime}(0)=0$ and $u(0)=0$, so that 0 is the minimum of $u$, which must then be non-negative.
For the second claim, fix any $\lambda\neq 0$. If $\alpha\neq 0$, then $E^{\mathrm{H}}_{\alpha}(\mu)=e^{-\alpha^{2}/8}<1=E_{\lambda}(\mu)$. If $\alpha=0$, then $E^{\mathrm{H}}_{\alpha}$ is identically equal to 111, and $\max(E_{\lambda}(0),E_{\lambda}(1))>1$, so we conclude.$\qed$
A.2 Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$